# One-Shot Method for Computing Generalized Winding Numbers

### v1

@@ 8 Conclusions
We presented a new method for computing a generalized winding number for parametric surfaces, meshes, and curve networks with minimal surfaces. With the use of SOS, our method can compute winding numbers within machine precision without surface discretization. We hope that the future work will explore applications of our method to tracking winding numbers of moving geometry, where discretization becomes even more of an issue.
Figure 9: Logarithmic error plot with respect to ground truth for our method (blue), Jacobson et al. (2013) (HGWN, red) and Barill et al. (2018) (FWN, green). Even when the boundary is discretized, our method computes winding numbers a few orders of magnitude more precisely than Jacobson et al. (2013) or Barill et al. (2018) on a mesh of comparable complexity. Horizontal axis shows the number of faces (top) and boundary vertices (bottom) of a mesh and the number of boundary vertices for our method.
@@ Acknowledgements
We gratefully acknowledge Ivan Puhachov, whose help with figures was invaluable, and to Paul Zhang for his assistance with Sum-of-squares Programming.
Figure 13: The results of our one-shot algorithm on meshes. Camel head mesh Sorkine (2005), Vase by virtox.
Figure 14: Additional results on parametric surfaces: a spiral with high winding numbers (a) and a non-manifold boat with Coons patches (b).
Figure 16: Curve networks can, depending on the context, represent different surfaces, such as (a) parametric patches Bessmeltsev et al. (2012), (b) minimal surfaces Wang and Chern (2021), Weber (2019) (c) other PDEs Nealen et al. (2007), or (d) more complex surfaces Yu et al. (2022).

### v2

@@ 8 Conclusions
We have presented a new method for computing a generalized winding number for parametric curves and surfaces, meshes, and curve networks with minimal surfaces. We show that for many cases such as 2D parametric curves or 3D meshes or parametric surfaces with simple boundary, on a regular grid, our method is faster than the state of the art. With the use of SOS, our method can compute winding numbers of parametric geometry precisely. We hope that the future work will explore applications of our method to tracking winding numbers of moving or optimized geometry, especially if the boundary remains fixed, where discretization becomes even more of an issue.
@@ Acknowledgements
We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC) under Grant No.: RGPIN-2019-05097 (“Creating Virtual Shapes via Intuitive Input”) and RGPIN-2024-04968 ("Modelling and animation via intuitive input"), and the NSERC - Fonds de recherche du Québec - Nature et technologies (FRQNT) NOVA Grant No. 314090. We thank Ivan Puhachov for providing several figures and Paul Zhang for his valuable assistance with Sum-of-Squares.
Appendix A Proof of Lemma 4.1
Intuitively, in $\mathbb{R}^{3}$, a ray and a surface intersect transversally if at each intersection point the ray is not tangent to the surface. At such intersection, the ray’s tangent and the surface’s tangent plane span the whole $\mathbb{R}^{3}$. More formally,
Definition A.1
[GP10]. The smooth map $f$ is said to be transversal to the submanifold $Z$, if at every point $x$ in the preimage of $Z$, where $f(x)=y$ and $T$ is a tangent space,
$\mathrm{Image}(df_{x})+T_{y}(Z)=T_{y}(Y).$
Figure 19: Two curves in 2D are transversal (a) if for every intersection their tangents at the intersection spans $\mathbb{R}^{2}$. A curve and a surface in 3D are transversal (c) if for every intersection the curve’s tangent and the surface’s tangent plane spans $\mathbb{R}^{3}$. Two submanifolds are not transversal if their tangent spaces do not span the ambient space (b,d).
In our context, the orientation number at an intersection point measures whether the ray and the normal to the surface are pointing in the same direction, i.e., it is $\mathrm{Sign}(n\cdot\overrightarrow{pq})$. For a more formal definition, please see [GP10].
In $Y=\mathbb{R}^{3}$, we first examine the case of a smooth surface $Z$ without boundary. The smooth map $f:X\rightarrow Y$ is a parameterization of a ray $f:\mathbb{R}^{+}\rightarrow\mathbb{R}^{3}$. More generally,
Proposition A.1
[GP10] If $f:X\rightarrow Y$ is transversal to $Z$, then $f^{-1}(Z)$ is a finite number of points, each with an orientation number $\pm 1$. We define the intersection number $I(f,Z)$ to be the sum of these orientation numbers.
The main paper refers to the intersection number as the ‘number of signed intersections’ as it is customary in computer graphics [FGC23b].
For boundaryless $X$ and $Z$, when at least one of them is compact, [GP10] formulate the following proposition:
Proposition A.2
[GP10] Homotopic maps have the same intersection numbers.
We will use this result to show that for a smoothly rotating ray, forming a homotopy, the intersection number can only change when the rotating ray passes a boundary of a surface. More generally, we formulate our result:
Proposition A.3
Let $X$ be an oriented manifold without boundary, and $M$ be a oriented submanifold with boundary of $Y$, $\text{dim}\;X+\text{dim}\;M=\text{dim}\;Y$. Let $f_{0},f_{1}$ be smooth maps $X\rightarrow Y$, also $F$ be a homotopy between them, i.e., $F:[0,1]\times X\rightarrow Y,F(0)=f_{0},F(1)=f_{1}$, and $f_{0},f_{1}$ intersect $M$ transversally and for each $s,F(s)$ intersects the boundary $\partial M$ not more than once. Then the oriented intersection number $I(F,M)$ only changes when F(s,)F(s,*) intersects the boundary $\partial M$, and only by increments of 1.
Proof A.4.
. We denote as $\overline{M}$ an arbitrary closure of $M$ such that the smooth manifold $Z=M\cup\overline{M}$ and $M\cap\overline{M}=\emptyset$. A transversal intersection on $Z$ may belong to $M$ or $\overline{M}$, thus, the intersection number $I(F,Z)=I(F,M)+I(F,\overline{M})$. By proposition A.2, $I(F,Z)$ does not change for those values of $s$.
Suppose that $F(s,X)$ intersects the boundary of $M$ only at $s=s_{0}$, maximum at one point. For this point, as the boundary is crossed to $s_{0}+\varepsilon$, the number of intersection points is no longer one, but still finite due to transversality, so $\overline{M}$ gains (or loses) this intersection, and since $I(F,Z)$ is constant, $I(F,M)$ loses (or gains) an intersection (Fig. 20). Therefore, the only value of $s$ where $I(F,M)$ can change is $s=s_{0}$ and the intersection number can only change by 1. Q.E.D.
Figure 20: The intersection number $I(F,Z)$ remains unchanged. When $Z$ is decomposed into a manifold $M$ with boundary and a closure $\overline{M}$ the intersection number $I(F,M)$ only changes at the boundary.
The only technicality is that in the propositions we assumed boundaryless $X$, while $\mathbb{R}^{+}$ has a boundary $\{0\}$. This is not an actual limitation: our homotopy $F(s,X)$ never changes the value $F(s,0)$, so the proposition still stands.
Appendix B Boundary Element Method Details
We follow the same notations and general setup as in Sec. 6.2.
For the flat 2D Laplace’s equation, the fundamental solution is
[…]
To compute the sign of an intersection, we look at the sign of the dot product between the normal of the surface at the intersection point, computed via differentiating Eq. (6), and the ray direction, e.g.:
$\frac{\partial f}{\partial u}=\int_{\eta\in\Gamma}\frac{\partial G(\xi,\eta)}{\partial u}\frac{\partial f(\eta)}{\partial n}-\frac{\partial^{2}G(\xi,\eta)}{\partial n\partial u}f(\eta)d\Gamma,$ (8)
where $\frac{\partial G(\xi,\eta)}{\partial\xi}$ and $\frac{\partial^{2}G(\xi,\eta)}{\partial n\partial\xi}$ are easily calculated in closed form (see Eq. 7).
Appendix C Sum-of-Squares Formulations
C.1 Intersections of Bézier curve projections
Here we outline the formulation for finding an intersection of two degree two Bézier curves, when projected onto a unit sphere via Sum-of-Squares (SOS) relaxation [MZPS21]. This formulation can be easily adapted to degree 3 Bézier curves with a simple change of basis functions. We are following the notations of [ZMST23].
In general, the problem of finding an intersection of spherical projections of two polynomial curves $f_{1}(t_{1}),f_{2}(t_{2}):[0,1]\rightarrow\mathbb{R}^{3}$ can be formulated as a following low-order polynomial optimization:
$\begin{split}&\min_{t_{1},t_{2},R}t_{1}^{2}\\ &f_{1}(t_{1})R=f_{2}(t_{2})\\ &t_{1},t_{2}\in[0,1],R>0\end{split}$ (9)
Concretely, for Bézier curves, basis functions are:
$\begin{split}&\phi_{1}(t):=(1-t)^{2}\\ &\phi_{2}(t):=2t(1-t)\\ &\phi_{3}(t):=t^{2}.\end{split}$ (10)
Then we can define two Bézier curves as $f_{1},f_{2}:[0,1]\rightarrow\mathbb{R}^{3}$, where $F_{i},G_{i}\in\mathbb{R}^{3}$ are the control vertices, i.e., $f_{1}(t_{1})=\sum_{i=1}^{3}F_{i}\phi_{i}(t_{1})$ and $f_{2}(t_{2})=\sum_{i=1}^{3}G_{i}\phi_{i}(t_{2})$. Then Eq. 9 has the following SOS relaxation:
$\begin{split}&\gamma*=\left\{\begin{array}[]{l}\max_{\gamma,s_{g},p_{h}}\gamma\\ \textrm{s.t.}\;t_{1}^{2}-\gamma-\sum_{g\in\mathcal{G}}s_{g}g-\sum_{h\in\mathcal{H}}p_{h}h\in\Sigma\\ s_{g}\in\Sigma_{d_{1}}\\ p_{h}\in\mathbb{R}[t]_{d_{2}}\\ \end{array}\right\}\\ &\mathcal{G}=\{(1-t_{1})t_{1},(1-t_{2})t_{2},R\}\\ &\mathcal{H}=f_{1}(t_{1})R-f_{2}(t_{2}),\end{split}$ (11)
The objective function $t_{1}^{2}$ is only needed to pick out a particular intersection. Once an intersection is found with values $t^{\prime}_{1},t^{\prime}_{2}$, we repeat the process after subdividing the curve at $t^{\prime}_{1}$ and taking its $t_{1}>t^{\prime}_{1}$ chunk.
Here $\Sigma_{d_{1}}$ are SOS polynomials of maximum degree $d_{1}$, and $\mathbb{R}[t]_{d_{2}}$ is a subset of polynomials of maximum degree $d_{2}$. This is a canonical SOS formulation, which can be solved by standard SDP solvers. In our implementation, we use YALMIP [Löf04, Löf09]. We use $d_{1}=4,d_{2}=2$ for both quadratic and cubic Bézier curves.
Self-intersections of a spherical projection of a Bézier curve are implemented in exactly the same way by substituting $f_{2}=f_{1}$ and replacing $t_{1}^{2}$ with $(t_{1}-t_{2})^{2}$. In this case, we ask to maximize the difference between $t_{1}$ and $t_{2}$. For this problem, solver does not indicate that a solution is not found if there is no intersection, so we need to check if the returned $t_{1},t_{2}$ are sufficiently different; in our implementation it is $|t_{1}-t_{2}|>10^{-6}$.
C.2 Ray Intersection with a Bézier Triangle
Similarly, we can consider the problem of finding the first intersection of a ray with a cubic Bézier triangle. As done in [MZPS21], this can be trivially generalized to other degrees of Bézier triangles or to an arbitrary NURBS patch.
A cubic Bézier triangle has 10 control vertices $F_{i}\in\mathbb{R}^{3}$ and is expressed as $f(u,v)=\sum_{i}\phi_{i}F_{i}$, where the basis functions $\phi_{i}$ are:
$\begin{split}&\phi_{1}(u,v)=-(u+v-1)^{3}\\ &\phi_{3}(u,v)=3u(u+v-1)^{2}\\ &\phi_{5}(u,v)=-6uv(u+v-1)\\ &\phi_{7}(u,v)=v^{3}\\ &\phi_{9}(u,v)=3u^{2}v\end{split}\begin{split}&\phi_{2}(u,v)=3v(u+v-1)^{2}\\ &\phi_{4}(u,v)=-3v^{2}(u+v-1)\\ &\phi_{6}(u,v)=-3u^{2}(u+v-1)\\ &\phi_{8}(u,v)=3uv^{2}\\ &\phi_{10}(u,v)=u^{3}.\end{split}$ (12)
The polynomial problem in Eq. (4) in the main document has the following SOS relaxation:
$\begin{split}&\gamma*=\left\{\begin{array}[]{l}\max_{\gamma,s_{g},p_{h}}\gamma\\ \textrm{s.t.}\;t^{2}-\gamma-\sum_{g\in\mathcal{G}}s_{g}g-\sum_{h\in\mathcal{H}}p_{h}h\in\Sigma\\ s_{g}\in\Sigma_{d_{1}}\\ p_{h}\in\mathbb{R}[t]_{d_{2}}\\ \end{array}\right\}\\ &\mathcal{G}=\{u,v,1-u,1-v,t\}\\ &\mathcal{H}=f(u,v)-O-tR.\end{split}$ (13)
We found that low degrees of $d_{1}=4,d_{2}=2$ are sufficient for exact recovery.